the perimeter of a semi-circle is doubled when the radius is increased by 6. Find the radius
Answer by dabanfield(803)

the perimeter of a semi-circle is doubled when the radius is increased by 6. Find the radius
Let R be the radius of the circle. Note that 2*R is equal to the diameter.
Then the perimeter of the semi-cirle is [(2*R)*pi)]/2 = R*pi.
If we increase R by 6 the perimeter is doubled so:
2*(R*pi) = (R+6)*pi
Solve for R.
Use pi = 3.1416.